# A Bayesian Estimator of Sample Size

### v1

@@ Abstract
We consider a Bayesian estimator of sample size (BESS) and an application to oncology dose optimization clinical trials. BESS is built upon balancing a trio of Sample size, Evidence from observed data, and Confidence in posterior inference. It uses a simple logic of “given the evidence from data, a specific sample size can achieve a degree of confidence in the posterior inference.” The key distinction between BESS and standard sample size estimation (SSE) is that SSE, typically based on Frequentist inference, specifies the true parameters values in its calculation while BESS assumes a possible outcome from the observed data. As a result, the calibration of the sample size is not based on Type I or Type II error rates, but on posterior probabilities. We argue that BESS leads to a more interpretable statement for investigators, and can easily accommodates prior information as well as sample size re-estimation. We explore its performance in comparison to SSE and demonstrate its usage through a case study of oncology optimization trial. BESS can be applied to general hypothesis tests. R functions are available at https://ccte.uchicago.edu/bess.
Keywords: Clinical Trial; Confidence; Evidence; Hypothesis; Posterior inference; Priors; Type I error.
@@ 1.1 Motivation
In novel drug development, the clinical objective is to establish the drug’s effectiveness and safety. Randomized clinical trials (RCTs) are the gold standard to achieve the objective. As it is usually impractical to include all patients from a disease population, RCTs take a random sample of certain size to address the clinical question through statistical inference. Due to the law of large numbers, the larger the sample size, the more precise the estimated treatment effect is but the more costly the trial is as well. Therefore, in practice a sample size estimation (SSE) approach is applied to compute an appropriate sample size for a trial to balance the tradeoff between precision in statistical inference and its cost.
We consider a Bayesian framework for SSE. The research is motivated by the recent change in early-phase oncology drug development that recommends randomized comparison of multiple doses. In the past two decades, oncology drug development has benefited from biological and genomics breakthroughs and novel cancer drugs are no longer based on cytotoxic one-size-fits-all mechanism like chemicals or radiations. Instead, targeted, immune, and gene and cell therapies combat tumor cells by precisely altering oncogenic cellular or molecular pathways. Consequently, the traditional monotonic dose-response relationship is no longer valid for many novel cancer drugs. Instead, efficacy often plateaus or even decreases after dose rises above certain level. To this end, US FDA launched Project Optimus (Food et al., 2023; Shah et al., 2021; Blumenthal et al., 2021) aiming to transform the early-phase oncology drug development. The main objective of the project is to identify an optimal dose, potentially lower than the maximum tolerated dose (MTD) but with at least comparable anti-tumor effect. The new dose optimization paradigm involves a randomized trial component comparing two or more doses following an initial dose escalation. While randomized trial design and SSE are routinely conducted in drug clinical trials, they are new in dose-optimization trials, usually coupled with limited resources. Traditional sample size estimation (SSE) is based on Frequentist Type I/II error rates, which is rarely the main objective in early-phase of drug development. Instead, investigators are more concerned about the accuracy in the trial decision making, such as selecting the right dose to start confirmatory studies which may be very costly. Current practice for SSE in dose optimization trials is often based on ad-hoc choices, such as using a small sample size of 20 patients per dose. This leads to a question of how much is expected to learn from the 20 patients, or whatever the number may be. Moreover, clinical trials that use human subjects for clinical research should always provide a reasonable justification for the number of subjects to be enrolled. We attempt to fill the gap by considering a Bayesian approach for sample size estimation.
@@ 1.2 Review of SSE Methods
In the literature, most SSE methods (Adcock, 1997; Wittes, 2002; Desu, 2012) are based on Frequentist hypothesis testing. A comprehensive review can be found in Wang and Ji (2020). Standard SSE methods determine the sample size by controlling the type I error rate $\alpha$ to achieve a desirable power $(1-\beta)$, assuming the true population parameters are known. For example, a sample size statement for a two-arm RCT based on binary outcome is as follows:
@@
In addition, sample size may be determined based on hybrid Frequentist-Bayesian inference as in Ciarleglio and Arendt (2017); Berry et al. (2010). These methods use Bayesian models but calibrate sample size based on type I error rate and power via simulation. In contrast, some methods use Bayesian properties for sample size estimation such as the length and coverage of posterior credible intervals (Pham-Gia et al., 1993; Joseph and Bélisle, 1997; M’Lan et al., 2008). Another type of Bayesian SSE methods determines sample size through Bayes factor with sampling and fitting priors (Lin et al., 2022), or loss functions like $k\cdot\text{False Positive Rate}+\text{False Negative Rate}$ (Müller et al., 2004).
Several works in the literature attempt to bridge the SSE philosophies between Frequentist and Bayesian methods. Notable contributions include Kunzmann et al. (2021), Inoue et al. (2005), and Lee and Zelen (2000). In particular, Lee and Zelen (2000) propose to link posterior estimates with Type I/II error rates, and estimate sample size by providing posterior probabilities rather than the Frequentist error rates. However, the sample size calculation still follows standard SSE approach. Notably the authors advocate defining “statistical significance” as the trial outcome having a large posterior probability of being correct.
@@ 1.3 Main idea
To this end, we propose a new Bayesian estimator of sample size (BESS) based on a general hierarchical modeling framework and posterior inference. Let $\bm{y}$ denote the observed data and $\bm{\theta}$ the unknown parameters as generic notation. We assume a Bayesian hierarchical model is to be used for data analysis. Our main idea is originated from noticing a simple trend between the sample size and observed effect size from data. Suppose in a two-arm RCT with sample size $n$ per arm and a binary outcome, we are interested in testing if the effect size (difference in response rates), defined as $\theta=(\theta_{1}-\theta_{0})$ between the treatment (1) and control (0) is greater than a clinically minimum effect size $\theta^{*}$. This can be formulated as testing the null hypothesis $H_{0}:\theta\leq\theta^{*}$ versus the alternative $H_{1}:\theta>\theta^{*}.$ Suppose the observed response rates are $\bar{y}_{1}$ and $\bar{y}_{0}$ for the treatment and control arms, respectively. A Frequentist $z$-test is used to test whether $\theta$ is greater than the minimum effect size $\theta^{*}$, given by
@@
Instead of using a Frequentist inference like the $z$-test, we consider Bayesian hypothesis testing based on posterior probability of the alternative hypothesis, which, in the setting of clinical trials, refers to the treatment being more effective than the control. Decision of accepting the alternative hypothesis is by thresholding the posterior probability of $H_{1}$ at a relatively large value $c\in(0,1)$. Assume a Bayesian hierarchical model is given by $f(\bm{y}\mid\bm{\theta})\pi(\bm{\theta}\mid H)\pi(H)$, where $H=H_{0}$ or $H_{1}$ is a binary indicator of the null and alternative hypotheses. The proposed BESS aims to find a balance between 1) “Sample size” $S$ of the clinical trial, 2) “Evidence” defined as the observed treatment effect $e=(\bar{y}_{1}-\bar{y}_{0})$, and 3) “Confidence” defined as the posterior probability of the alternative hypothesis.
@@
Note that “confidence” here refers to $\text{Pr}(H_{1}|\bm{y})$, the posterior probability of the alternative hypothesis, i.e., the treatment effect is at least $\theta^{*}.$ Comparing Statement 1 and Statement 2, one can see that the two statements are based on different statistical properties, Frequentist type I/II error rates for Statement 1 and Bayesian posterior probabilities for Statement 2. We will show that the BESS and associated Statement 2 is easier to interpret in practice since it directly addresses the uncertainty in the decision to be made for the trial at hand, measured by posterior probabilities.
@@ 2 Probability Model
We consider BESS for both one-arm and two-arm trials. Denote $y_{ij}$ the outcome of patient $i$ in arm $j$, where $i=1,\ldots,n,$ index the patients, and $j=0$ and $1$ index the control and treatment arm, respectively. When needed, we drop index $j$ and use $y_{i}$ for one-arm trials. Let $\theta_{1}$ denote the true response parameter for the treatment arm, and let $\theta_{0}$ be the true response parameter for the control arm in a two-arm trial or the reference response parameter in a one-arm trial. Let $\theta=(\theta_{1}-\theta_{0})$ be the true treatment effect. Consider hypotheses
@@
where $f(.)$ represents the likelihood function, $\pi(\theta_{0},\theta_{1}|\tilde{\bm{\theta}})$ is a joint prior distribution for $(\theta_{0},\theta_{1})$, $\tilde{\bm{\theta}}$ are hyper-parameters, and $I(x\in A)$ is the indicator function which equals 1 if $x\in A$ and 0 otherwise. For simplicity, we consider $\pi(\theta_{0},\theta_{1}|\tilde{\bm{\theta}})=\pi_{0}(\theta_{0}|\tilde{\bm{\theta}})\pi_{1}(\theta_{0}|\tilde{\bm{\theta}})$, where $\pi_{0}=\pi_{1}$. In this work, we consider three specific types of outcome $y_{ij}$: binary, continuous, and count. A summary of the parameter $\theta_{j}$, likelihood function $f(.)$, and prior distribution $\pi_{j}$ for $\theta_{j}$ is shown in Table 1. In all three cases, the observed sample mean for each arm $j$, denoted as $\bar{y}_{j}$, is the sufficient statistics for $\theta_{j}$. In Table 1 we consider conjugate prior for simplicity and computational speed. In general, different priors can be considered. While it is not the focus of this work to discuss choice of priors, we note the flexibility of BESS to incorporate various priors in real-life applications. For example, when little prior information is known vague priors like Jeffrey’s prior may be considered; in contrast, it is also possible to use informative priors when prior information is available. For instance, assume a previous trial with binary outcome has completed with a sample size of $n_{0}$ patients of whom their outcome data are available, denoted as $\bm{y}^{0}=\{y_{i}^{0};i=1,\ldots,n_{0}\}$. Then, one could consider an informative prior $\pi_{1}(\theta_{1})=\text{Beta}(a,b)$ and set $a=a^{*}+\sum_{i=1}^{n_{0}}y_{i}^{0},\,\,b=b^{*}+n_{0}-\sum_{i=1}^{n_{0}}y_{i}^{0},$ where $a^{*}$ and $b^{*}$ are small (e.g., $a^{*}=b^{*}=0.5$).
@@ 3.1 Confidence
We introduce the three pillar of BESS, Confidence, Evidence, and Sample Size. We start with “C”, the confidence, which refers to the confidence of posterior inference expressed mathematically as the posterior probability of the alternative hypothesis $\text{Pr}(H=H_{1}|\bm{y}_{n}),$ where $\bm{y}_{n}$ denotes the data with sample size $n$. The optimal decision rule under a variety of loss functions (Müller et al., 2004) is to reject the null $H_{0}$ and accept $H_{1}$ if $\text{Pr}(H=H_{1}|\bm{y}_{n})\geq c$ for cutoff $c\in(0,1).$ The value of $c$ measures the least confidence of the decision to reject $H_{0}$ and accept $H_{1}$. The higher $c$ is, the more confident is the decision. To see this, one simply observes that $(1-c)$ is the upper bound of posterior probability of a wrong rejection since when $H_{1}$ is rejected, $\text{Pr}(H=H_{0}|\bm{y}_{n})<c$.
@@ 3.2 Evidence
In simple words, evidence $e$ is the observed effect size from the trial data before they are observed. This means that in order to apply BESS, investigator needs to pre-specify (and calibrate) the potential observed effect size before the trial is conducted. This is analogous to the requirement of specifying the true parameter values in standard SSE, except BESS assumes what might be observed rather than what might be true. We consider evidence $e$ as a function of the sufficient statistic in Table 1. For the three outcomes in one-arm trials, $(e+\theta_{0})$ is exactly the sufficient statistic. For two-arm trials, $e$ is the difference between the sufficient statistics $(\bar{y}_{0},\bar{y}_{1})$ of the two arms. Being a function of sufficient statistics allows for a search algorithm to find an appropriate sample size under BESS, which will be clear next.
@@ 3.3 Sample Size of BESS
We first briefly review the standard SSE based on Frequentist inference. Consider a $z$-test for a two-arm trial with binary outcome, the standard sample size approach assumes true values of $\theta_{1}$ and $\theta_{0}$, and solves for $n$ based on desirable Type I/II error rates $\alpha$/$\beta$ given by
@@
BESS 2.1: Assuming the evidence is $e$, $n$ subjects per arm are needed to declare with confidence $c$ that the treatment effect is larger than $\theta^{*}$, assuming a known variance of $\sigma^{2}$ in each arm.
@@
Input: The hierarchical models in Table 1.
Input: Clinically meaningful effect size $\theta^{*}$, evidence $e$, confidence $c$, prior probability $q$, reference response rate $\theta_{0}$ (for one-arm trials).
Set $n_{\text{min}}$ and $n_{\text{max}}$ ($n_{\text{min}}<n_{\text{max}}$) the smallest and largest candidate sample sizes.
Set $n=n_{\text{min}}.$
while $n\leq n_{\text{max}}$ do
     Compute equation (3).
     if condition (7) is true then
         Stop and return the sample size $n$.
     else
         $n=n+1$
     end if
end while
if $n>n_{\text{max}}$ then
     Return sample size is larger than $n_{\text{max}}.$
end if
@@
For binary and count data in two-arm trials, $\text{Pr}(H=H_{1}|\bm{y}_{n})=\text{Pr}(H=H_{1}|\bar{y}_{1},\bar{y}_{0},n)$. See Appendix A.3 for detail. Since $e=\bar{y}_{1}-\bar{y}_{0}$, fixing evidence $e$ does not uniquely define $\text{Pr}(H=H_{1}|\bm{y}_{n}).$ Hence, we propose to specify evidence $e$ and find all pairs of $\bar{y}_{1}$ and $\bar{y}_{0}$ that satisfies $\bar{y}_{1}-\bar{y}_{0}=e$. We then find the minimum posterior probability of $H_{1}$ among these pairs of $\bar{y}_{1}$ and $\bar{y}_{0}$, i.e.,
@@
We propose BESS Algorithm 2 and the corresponding sample size statement BESS 2.2.
BESS 2.2 Assuming the evidence is $e$, at least $n$ subjects per arm are needed to declare with confidence $c$ that the treatment effect is larger than $\theta^{*}$.
@@
Alternatively, one may wish to specify the values of $\bar{y}_{1}$ and $\bar{y}_{0}$ directly instead of their difference $e$. Since $(\bar{y}_{1},\bar{y}_{0})$ is sufficient, the sample size search algorithm becomes easier. To this end, we propose a simple variation BESS Algorithm 2’ in Appendix A.4 and statement BESS 2.2’ assuming $(\bar{y}_{1},\bar{y}_{0})$ are given.
BESS 2.2’ Assuming the response parameters in treatment and control arms are $\bar{y}_{1}$ and $\bar{y}_{0}$, respectively, $n$ subjects per arm are needed to declare with confidence $c$ that the treatment effect is larger than $\theta^{*}$.

### v2

@@ Abstract
We consider a Bayesian estimator of sample size (BESS) and an application to oncology dose optimization clinical trials. BESS is built upon three pillars, Sample size, Evidence from observed data, and Confidence in posterior inference. It uses a simple logic of “given the evidence from data, a specific sample size can achieve a degree of confidence in the posterior inference.” The key distinction between BESS and standard sample size estimation (SSE) is that SSE, typically based on Frequentist inference, specifies the true parameters values in its calculation while BESS assumes possible outcome from the observed data. As a result, the calibration of the sample size is not based on type I or type II error rates, but on posterior probabilities. We demonstrate that BESS leads to a more interpretable statement for investigators, and can easily accommodates prior information as well as sample size re-estimation. We explore its performance in comparison to the standard SSE and demonstrate its usage through a case study of oncology optimization trial. BESS can be applied to general hypothesis tests. An R tool is available at https://ccte.uchicago.edu/BESS.
Clinical Trial; Confidence; Evidence; Hypothesis testing; Posterior inference; Priors; Type I error.
@@ 1.1 Motivation
In drug development, the clinical objective is to establish the drug’s effectiveness and safety. Randomized clinical trials (RCTs) are the gold standard to achieve the objective. As it is usually impractical to include all patients from a disease population, RCTs take a random sample of certain size to address the clinical question through statistical inference. Due to the law of large numbers, the larger the sample size, the more precise the estimated treatment effect is but also the more costly the trial is. Therefore, in practice a sample size estimation (SSE) is usually conducted before the trial starts to balance the tradeoff between precision in statistical inference and trial cost.
We consider a Bayesian framework for SSE. The research is motivated by the recent change in early-phase oncology drug development that recommends randomized comparison of multiple doses. In the past two decades, oncology drug development has benefited from biological and genomics breakthroughs and novel cancer drugs are no longer based on cytotoxic one-size-fits-all mechanism like chemicals or radiations. Instead, targeted, immune, and gene and cell therapies combat tumor cells by precisely altering oncogenic cellular or molecular pathways. Consequently, the traditional monotonic dose-response relationship is no longer valid for many novel cancer drugs. Instead, efficacy often plateaus or even decreases after dose rises above certain level. To this end, US FDA launched Project Optimus (Food et al., 2023; Shah et al., 2021; Blumenthal et al., 2021) aiming to transform the early-phase oncology drug development. The main objective of the project is to identify an optimal dose, potentially lower than the maximum tolerated dose but with at least comparable anti-tumor effect. The new dose optimization paradigm recommends a randomized comparison of two or more doses following an initial dose escalation. While randomized trial design and SSE are routinely conducted in drug clinical trials, they are new in dose-optimization trials, usually restricted by limited resources. Type I/II error rates for standard SSE are rarely the main objective for early-phase of drug development. Instead, investigators are more concerned about the accuracy in the trial decision making, such as selecting the right dose or starting confirmatory studies, which may be very costly. Current SSE for dose optimization trials is often based on ad-hoc choices, such as using a small sample size of 20 patients per dose. This leads to a question of how much is expected to learn from the 20 patients, or whatever the number may be. Moreover, clinical trials that use human subjects for clinical research should always provide a reasonable justification for the number of subjects to be enrolled. We attempt to fill the gap by considering a Bayesian approach for sample size estimation.
@@ 1.2 Review of SSE Methods
In the literature, standard SSE methods (Adcock, 1997; Wittes, 2002; Desu, 2012) are based on Frequentist hypothesis testing by controlling the type I error rate $\alpha$ to achieve a desirable power $(1-\beta)$, assuming the true population parameters are known. For example, a sample size statement for a two-arm RCT based on binary outcome is as follows:
@@
A comprehensive review can be found in Wang and Ji (2020). In addition, sample size may be determined based on hybrid Frequentist-Bayesian inference as in Ciarleglio and Arendt (2017); Berry et al. (2010). These methods use Bayesian models but calibrate sample size based on type I error rate and power via simulation. In contrast, some methods use Bayesian properties for sample size estimation such as the length, coverage of posterior credible intervals, Bayes factors under sampling and fitting priors (Lin et al., 2022), or loss functions like $k\cdot\text{False Positive Rate}+\text{False Negative Rate}$ (Müller et al., 2004).
Several methods attempt to bridge the SSE philosophies between Frequentist and Bayesian methods. Notable contributions include Kunzmann et al. (2021), Inoue et al. (2005), and Lee and Zelen (2000). In particular, Lee and Zelen (2000) propose to link posterior estimates with type I/II error rates, and estimate sample size by providing posterior probabilities rather than the Frequentist error rates. The authors advocate defining “statistical significance” as the trial outcome having a large posterior probability of being correct. However, their sample size calculation still follows standard SSE approach.
@@ 1.3 Main idea
To this end, we propose a new Bayesian estimator of sample size (BESS) based on a general hierarchical modeling framework and posterior inference. Let $\bm{y}$ denote the observed data and $\bm{\theta}$ the unknown parameters as generic notation. We assume a Bayesian hierarchical model is to be used for the experimental design and data analysis. Our main idea is originated from noticing a simple trend between the sample size and observed effect size from data. Suppose in a two-arm RCT with sample size $n$ per arm and a binary outcome, we are interested in testing if the effect size (difference in response rates), defined as $\theta=(\theta_{1}-\theta_{0})$ between the treatment (1) and control (0) is greater than a clinically minimum effect size $\theta^{*}$. This can be formulated as testing the null hypothesis $H_{0}:\theta\leq\theta^{*}$ versus the alternative $H_{1}:\theta>\theta^{*}.$ Suppose after the trial is completed, the observed response rates are $\bar{y}_{1}$ and $\bar{y}_{0}$ for the treatment and control arms, respectively. A Frequentist $z$-test is used to test the hypotheses given by
@@
[…]
Instead of using a Frequentist inference like the $z$-test, we consider Bayesian hypothesis testing based on posterior probability of the alternative hypothesis, which, in the setting of clinical trials, refers to the success of the trial. Decision of accepting the alternative hypothesis is by thresholding the posterior probability of $H_{1}$ at a relatively large value $c\in(0,1)$. Assume a Bayesian hierarchical model is given by $f(\bm{y}\mid\bm{\theta})\pi(\bm{\theta}\mid H)\pi(H)$, where $H=H_{0}$ or $H_{1}$ is a binary indicator of the null and alternative hypotheses. The proposed BESS aims to find a balance between 1) “Sample size” $S$ of the clinical trial, 2) “Evidence” defined as the observed treatment effect $e=(\bar{y}_{1}-\bar{y}_{0})$, and 3) “Confidence” defined as the posterior probability of the alternative hypothesis.
@@
Note that “evidence” is a function of the data, not parameters, and “confidence” refers to $\text{Pr}(H_{1}|\bm{y})$, the posterior probability of the alternative hypothesis, i.e., the treatment effect is at least $\theta^{*}.$ Comparing Statement 1 and Statement 2, one can see that the two statements are based on different statistical properties, Frequentist type I/II error rates for Statement 1 and Bayesian posterior probabilities for Statement 2. Also, while Statement 1 assumes the true values of the parameters are given, Statement 2 assumes what might be observed from the trial data. We will show that the BESS and associated Statement 2 is easier to interpret in practice since it directly addresses the uncertainty in the decision to be made for the trial at hand, measured by posterior probabilities.
@@ 2 Probability Model
We consider BESS for both one-arm and two-arm trials. Denote $y_{ij}$ the outcome of patient $i$ in arm $j$, where $i=1,\ldots,n,$ index the patients, and $j=0$ and $1$ index the control and treatment arm, respectively. When needed, we drop index $j$ and use $y_{i}$ for one-arm trials. Let $\theta_{1}$ denote the true response parameter for the treatment arm, and let $\theta_{0}$ be the true response parameter for the control arm in a two-arm trial or the reference response parameter in a one-arm trial. Let $\theta=f(\theta_{0},\theta_{1})$ a function of $\theta_{0}$ and $\theta_{1}$ denote the treatment effect. For example, we consider $\theta=(\theta_{1}-\theta_{0})$ in this paper although its form can be generalized. Consider hypotheses
@@
where $f(.)$ represents the likelihood function, $\pi(\theta_{0},\theta_{1}|\tilde{\bm{\theta}})$ is a joint prior distribution for $(\theta_{0},\theta_{1})$, $\tilde{\bm{\theta}}$ are hyper-parameters, and $I(x\in A)$ is the indicator function which equals 1 if $x\in A$ and 0 otherwise. For simplicity, we consider $\pi(\theta_{0},\theta_{1}|\tilde{\bm{\theta}})=\pi_{0}(\theta_{0}|\tilde{\bm{\theta}})\pi_{1}(\theta_{0}|\tilde{\bm{\theta}})$, where $\pi_{0}=\pi_{1}$.
In this work, we consider three specific types of outcome $y_{ij}$: binary, continuous, and count. A summary of the parameter $\theta_{j}$, likelihood function $f(.)$, and prior distribution $\pi_{j}$ for $\theta_{j}$ is shown in Table 1. We consider conjugate prior for simplicity and computational convenience although the proposed BESS works for any general priors. While it is not the focus of this work to discuss choice of priors, we note the flexibility of BESS to incorporate various priors in real-life applications. For example, when little prior information is known vague priors like Jeffrey’s prior may be considered; in contrast, it is also possible to use informative priors when prior information is available. For instance, assume a previous trial with binary outcome has completed with a sample size of $n_{0}$ patients of whom their outcome data are available, denoted as $\bm{y}^{0}=\{y_{i}^{0};i=1,\ldots,n_{0}\}$. Then, one could consider an informative prior $\pi_{1}(\theta_{1})=\text{Beta}(a,b)$ and set $a=a^{*}+\sum_{i=1}^{n_{0}}y_{i}^{0},\,\,b=b^{*}+n_{0}-\sum_{i=1}^{n_{0}}y_{i}^{0},$ where $a^{*}$ and $b^{*}$ are small (e.g., $a^{*}=b^{*}=0.5$).
@@ 3.1 Confidence
We introduce the three pillars of BESS, Confidence, Evidence, and Sample Size. We start with “Confidence”, the confidence of posterior inference, expressed mathematically as the posterior probability of the alternative hypothesis $\text{Pr}(H=H_{1}|\bm{y}_{n}),$ where $\bm{y}_{n}$ denotes the data with sample size $n$. The optimal decision rule under a variety of loss functions (Müller et al., 2004) is to reject the null $H_{0}$ and accept $H_{1}$ if $\text{Pr}(H=H_{1}|\bm{y}_{n})\geq c$ for a high value of $c\in(0,1).$ The higher $c$ is, the more confident is the decision. To see this, one simply observes that $(1-c)$ is the upper bound of posterior probability of a wrong rejection since when $H_{0}$ is rejected, $\text{Pr}(H=H_{0}|\bm{y}_{n})<c$.
@@ 3.2 Evidence
In simple words, evidence $e$ is the observed effect size from the trial data before they are observed. This means that in order to apply BESS, investigator needs to pre-specify (and calibrate) the potential observed effect size before the trial is conducted. This is analogous to the requirement of specifying the true parameter values in the standard SSE, except BESS assumes what might be observed rather than what might be true. We consider evidence $e$ as a function of the sufficient statistic in Table 1. For the three outcomes in one-arm trials, $(e+\theta_{0})$ is exactly the sufficient statistic. For two-arm trials, $e$ is the difference between the sufficient statistics $(\bar{y}_{0},\bar{y}_{1})$ of the two arms. Being a function of sufficient statistics allows for a search algorithm to find an appropriate sample size under BESS, which will be clear next.
@@ 3.3 Sample Size of BESS
We first briefly review the standard SSE based on Frequentist inference. Considering a $z$-test for a two-arm trial with binary outcome, the standard sample size approach assumes true values of $\theta_{1}$ and $\theta_{0}$, and solves for $n$ based on desirable type I/II error rates $\alpha$/$\beta$ given by
@@
BESS 2.1: Assuming the evidence is $e$ and variance $\sigma^{2}$ is known for each arm, $n$ subjects per arm are needed to declare with confidence $c$ that the treatment effect is larger than $\theta^{*}$.
BESS Algorithm 1 below provides steps for finding $n$ in BESS 1 and BESS 2.1.
@@
1:Input: The hierarchical models in Table 1.
2:Input: Clinically meaningful effect size $\theta^{*}$, evidence $e$, confidence $c$, prior probability $q$, reference response rate $\theta_{0}$ (for one-arm trials).
4:Set $n=n_{\text{min}}.$
5:while $n\leq n_{\text{max}}$ do
6:     Compute equation (3).
7:     if condition (7) is true then
8:         Stop and return the sample size $n$.
9:     else
10:         $n=n+1$
11:     end if
12:end while
13:if $n>n_{\text{max}}$ then
14:     Return sample size is larger than $n_{\text{max}}.$
15:end if
@@
For binary and count data in two-arm trials, $\text{Pr}(H=H_{1}|\bm{y}_{n})=\text{Pr}(H=H_{1}|\bar{y}_{1},\bar{y}_{0},n)$. See Appendix A.3 for a quick proof. Since $e=(\bar{y}_{1}-\bar{y}_{0})$, fixing evidence $e$ does not uniquely define $\text{Pr}(H=H_{1}|\bm{y}_{n}).$ Hence, we propose to specify evidence $e$ and find all pairs of $\bar{y}_{1}$ and $\bar{y}_{0}$ that satisfies $e=(\bar{y}_{1}-\bar{y}_{0})$. We then find the minimum posterior probability of $H_{1}$ among these pairs of $\bar{y}_{1}$ and $\bar{y}_{0}$, i.e.,
@@
Then the sample size statement for a two-arm trial is the same as BESS 1. And BESS Algorithm 2 provides details of finding the sample size $n$ for a two-arm trial.
@@
Alternatively, one may wish to specify the values of $\bar{y}_{1}$ and $\bar{y}_{0}$ directly instead of their difference $e$. Since $(\bar{y}_{1},\bar{y}_{0})$ is sufficient, the sample size search algorithm becomes easier. To this end, we propose a simple variation BESS Algorithm 2’ in Appendix A.4 and statement BESS 2.2 assuming $(\bar{y}_{1},\bar{y}_{0})$ are given.
BESS 2.2 Assuming the response parameters in treatment and control arms are $\bar{y}_{1}$ and $\bar{y}_{0}$, respectively, $n$ subjects per arm are needed to declare with confidence $c$ that the treatment effect is larger than $\theta^{*}$.